# Segment Anything Model (SAM) for Radiation Oncology

### v1

@@ 5.2.3 Collaborative AI-Human Decision Making
SAM can act as a decision support system, not just providing initial segmentation but also flagging complex cases that require more in-depth scrutiny by the clinician. This dual workflow could optimize the clinician’s efforts towards cases that are more critical, improving overall treatment efficacy.
5.3 Application Cases of SAM
5.3.1 Automatic Delineation of OARs
SAM’s use can be expanded towards challenging tasks like delineating OARs with similar densities to the target region, such as the rectum in prostate cancer cases, or parotid glands in head and neck cancers. With enough training on these specific regions, SAM can become a crucial tool in minimizing radiation-induced complications [82, 83].
@@ 5.3.2 Inter-Patient Adaptation
SAM can also be tailored to aid inter-patient adaptation. Its ability to automatically adjust segmentation masks from one patient to another, factoring in individual anatomy and pathology, can be utilized to predict how a tumor might evolve or migrate over time. It could thus help in adapting the treatment plan for more personalized and efficient radiation therapy.
5.4 Improving SAM for Radiation Oncology
@@ 5.4.1 Custom Training for Specific Tasks
A promising method to enhance SAM’s performance for complex organs involves task-specific fine-tuning. This could be realized by training SAM with an enriched dataset, containing a wide variety of examples of these complex structures.
@@
Another critical feature of these adapters is their ability to operate in 3D space, providing the capability to process 3D medical scans, a common data modality in radiation oncology [85]. This would enable the model to better capture the spatial and depth correlations within the data, a crucial factor for accurate delineation of complex structures.
The adapter modules can be trained using a variety of self-supervised learning methods, such as contrastive embedding-mixup and shuffled embedding prediction [86, 84, 87], thereby enhancing their ability to handle complex segmentation tasks. These techniques may provide a robust solution to improve SAM’s performance on complex organ segmentation, thereby potentially revolutionizing its utility in radiation oncology.
@@ 5.4.2 Enhancement with Language Prompts and LLM Encoder
Another promising direction for enhancing SAM involves integrating it with language prompts and language model encoder [88, 89, 90]. Language prompts can be leveraged to guide the segmentation process, providing context or highlighting areas of interest in the image. In such a setup, the LLM encoder can convert the language prompts into a meaningful context vector that guides SAM’s segmentation process. By aligning the language and vision model training through a shared context representation [88], SAM may better delineate complex structures and ambiguous regions. This approach capitalizes on the strength of LLMs to understand and generate nuanced human language, potentially enabling more precise and clinically relevant segmentations.
@@ 5.4.3 Reinforcement Learning and Expert Feedback
In reinforcement learning paradigms [91, 92], the system learns by interacting with its environment (in this case, the imaging data), and improves based on feedback (the differences between SAM’s segmentations and expert contours). SAM’s "actions," which in this context are the segmentations it produces, would be continuously adjusted based on "rewards," or the degree of similarity to expert-produced contours. This could lead to a system where SAM progressively adapts its segmentation approach to match expert standards more closely, with its progress validated against the expert contouring data.
@@ 5.4.4 Uncertainty Estimation
Implementing a measure of uncertainty in SAM’s output could be achieved by integrating Bayesian approaches into the learning process, such as Bayesian Convolutional Neural Networks [93]. These provide an estimate of the model’s confidence in its predictions, allowing clinicians to identify areas where manual review might be necessary.
@@ 5.4.5 Integration of Clinical Knowledge
SAM could be improved by integrating clinical information, such as the relationship between tumor location, size, and the expected effect of radiation. Machine learning models such as graph neural networks [94] could be used to encode this information, where nodes could represent different anatomical regions, and edges could represent spatial relationships or common patterns of metastasis. Combining this with SAM’s powerful image segmentation capabilities could yield a model that is both accurate and clinically relevant.
@@ 6 Conclusion
In this study, we explored the application of the Segment Anything Model (SAM) for medical image segmentation in radiation oncology. We found that SAM shows promising potential for aiding treatment planning, but it requires significant customization and refinement for clinical use. Our evaluation highlighted SAM’s capacity to delineate large, distinct organs effectively, but revealed difficulties in segmenting smaller, intricate structures, especially with ambiguous prompts. This variability in performance across different anatomical sites and imaging modalities underlines the necessity to tailor SAM and similar foundation models to the specific requirements of medical domains.

### v2

@@ 5.2.3 Collaborative AI-Human Decision Making
SAM can act as a decision support system, not just providing initial segmentation but also flagging complex cases that require more in-depth scrutiny by the clinicians. This dual workflow could optimize the clinician’s efforts towards cases that are more critical, improving overall treatment efficacy.
5.2.4 Automatic Segmentation of OARs
SAM’s use can be expanded towards challenging tasks like delineating OARs with similar densities as the targets, such as the rectum in prostate cancer cases, or parotid glands in head & neck cancers. With enough training on these specific sites, SAM can become a crucial tool in minimizing radiation-induced complications [98, 99].
@@ 5.2.5 Inter-Patient Adaptation
SAM can also be tailored to aid inter-patient adaptation. Its ability to automatically adjust segmentation masks from one patient to another, factoring in individual anatomy and pathology, can be utilized to predict how a tumor might evolve or migrate over time. It could thus help in adapting the treatment plan for more personalized and precision radiation therapy.
@@ 5.2.6 Custom Training for Specific Tasks
A promising method to enhance SAM’s performance for complex OARs involves task-specific fine-tuning. This could be realized by training SAM with an enriched dataset, containing a wide variety of examples of these complex structures.
@@
Another critical feature of these adapters is their ability to operate in 3D space, providing the capability to process 3D medical images, a common data format in radiation oncology [101]. This would enable the model to better capture the spatial and depth correlations within the data, a crucial factor for accurate delineation of complex structures.
The adapter modules can be trained using a variety of self-supervised learning methods, such as contrastive embedding-mixup and shuffled embedding prediction [102, 100, 103], thereby enhancing their ability to handle complex segmentation tasks. These techniques may provide a robust solution to improve SAM’s performance on complex OAR segmentation, thereby potentially revolutionizing its utility in radiation oncology.
@@ 5.2.7 Enhancement with Language Prompts and LLM Encoder
Another promising direction for enhancing SAM involves integrating it with language prompts and language models [23, 83, 104], especially in application scenarios where domain-specific language and knowledge are beneficial [105, 16, 106, 107, 108, 109]. Language prompts can be leveraged to guide the segmentation process, providing context or highlighting areas of interest in the images. In such a setup, the LLM encoder can convert the language prompts into a meaningful context vector that guides SAM’s segmentation process. By aligning the language and vision model training through a shared context representation [23], SAM may better delineate complex structures and ambiguous regions. This approach capitalizes on the strength of LLMs to understand and generate nuanced human language, potentially enabling more precise and clinically relevant segmentation.
@@ 5.2.8 Reinforcement Learning and Expert Feedback
In reinforcement learning paradigms [110, 111], the system learns by interacting with its environment (in this case, the imaging data), and improves based on feedback (the differences between SAM’s segmentation and expert-produced contours). SAM’s "actions," which in this context are the segmentation it produces, would be continuously adjusted based on "rewards," or the degree of similarity to expert-produced contours. This could lead to a system where SAM progressively adapts its segmentation approach to match expert-produced contours more closely, with its progress validated against the expert-produced contours.
@@ 5.2.9 Uncertainty Estimation
Implementing a measure of uncertainty in SAM’s output could be achieved by integrating Bayesian approaches into the learning process, such as Bayesian Convolutional Neural Networks [112]. These provide an estimate of the model’s confidence in its predictions, allowing clinicians to identify areas where manual review and edits might be necessary.
@@ 5.2.10 Integration of Clinical Knowledge
SAM could be improved by integrating clinical information, such as the relationship between tumor location, size, and the expected effect of radiation. Machine learning models such as graph neural networks [113] could be used to encode this information, where nodes could represent different anatomical sites, and edges could represent spatial relationships or common patterns of metastasis. Combining this with SAM’s powerful image segmentation capabilities could yield a model that is both accurate and clinically relevant.
@@ 6 Conclusion
In this study, we explored the application of the SAM for medical image segmentation in radiation oncology. We found that SAM shows promising potential for aiding treatment planning, but it requires significant customization and refinement for clinical use. Our evaluation highlighted SAM’s capacity to delineate large, distinct organs effectively, but revealed difficulties in segmenting smaller, intricate structures, especially with ambiguous prompts. This variability in performance across different anatomical sites and imaging modalities underlines the necessity to tailor SAM and similar foundation models to the specific requirements in medical domains.